Thompson and Thompson is a steel bolts manufacturing company. Their current steel bolts have a mean diameter of 147 millimeters, and a standard deviation of 6 millimeters. If a random sample of 33 steel bolts is selected, what is the probability that the sample mean would differ from the population mean by greater than 1.1 millimeters? Round your answer to four decimal places.
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 147
std deviation   =σ= 6.0000
sample size       =n= 33
std error=σ=σ/√n= 1.0445

: probability that the sample mean would differ from the population mean by greater than 1.1 millimeters :

probability = 1-P(145.9<X<148.1) = 1-P(-1.05<Z<1.05)= 1-(0.8531-0.1469)= 1-0.7062=0.2938